Find the nature of the planar section of the conicoid x^2÷3 −y^2÷4 = z by the plane
x+2y−z = 6
1
Expert's answer
2020-02-18T05:49:07-0500

We have system of equations (conicoid and plane):"\begin{cases}
\frac{x^2}{3}-\frac{y^2}{4}=z \\
x+2y-z=6
\end{cases}
\quad 
\begin{cases}
\frac{x^2}{3}-\frac{y^2}{4}=z\\
z=x+2y-6
\end{cases}
\quad 
\begin{cases}
\frac{x^2}{3}-\frac{y^2}{4}=x+2y-6\\
z=x+2y-6
\end{cases}"


"\frac{x^2}{3}-\frac{y^2}{4}=x+2y-6, \quad (\frac{x^2}{3}-x+\frac{3}{4})-\frac{3}{4}-(\frac{y^2}{4}+2y+4)+4=-6"


"(\frac{x}{\sqrt{3}}-\frac{\sqrt{3}}{2})^2-(\frac{y}{2}+2)^2=-9.25" - equation of the hyperbolic cylinder.


"\begin{cases}

 (\frac{x}{\sqrt{3}}-\frac{\sqrt{3}}{2})^2-(\frac{y}{2}+2)^2=-9.25
\\
z=x+2y-6
\end{cases}"


Intersection of hyperbolic cylinder and plane can be:

• a line, 2 parallel lines, no points (if plane is perpendicular to "z=0" )

• hyperbola (if if plane isn’t perpendicular to "z=0" )


In our case, "x+2y-z=6" isn’t perpendicular to "z=0".

Therefore, it will be hyperbola.


Answer: hyperbola.